# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v3

@@ 8.3 Sim-to-Real Gap Solutions
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (Zhang et al. (2024b)) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (Jiang et al. (2024a)) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
@@ 8.4 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (Mandlekar et al. (2023)) and DexMimicGen (Jiang et al. (2024b)) change the positions and orientations of objects, while CACTI (Mandi et al. (2022)) adds new, artificial objects to the scene. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Automated Generation methods. In the Real-to-Sim methods, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (Chen et al. (2024c)). Additionally, GRS (Zook et al. (2024)) utilizes VLM to extract 3D object meshes corresponding to real-world object from assert database based on RGB image. In the Automated Generation methods, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (Wang et al. (2023d)). When the output is scene code, it directly generates the corresponding scene (Wang et al. (2023b)). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (Wang et al. (2023d)). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@ 9 Discussion
In this survey, we aim to outline the opportunities brought by foundation models for general manipulation. We believe the potential of embedding foundation models into manipulation tasks as a viable path towards achieving general manipulation. However, the primary applications of LLMs, VFMs, VLMs, LMMs and VGMs focus only on certain aspects of general manipulation capability, such as reasoning, perception, multimodal understanding, and data generation. The current framework for RFMs resembles human learning, requiring no calibration or robotic parameters, directly learning the mapping between observations and action. However, this demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring over a 99% success rate remains an unresolved concern. Therefore, this paper proposes a framework of robot learning for manipulation towards achieving general manipulation capability and detailing how foundation models can address challenges in each module of the framework. However, there are still many open questions in this survey. In this section, we delve into several open questions that we are particularly concerned about.
@@ 9.3 What kind of learning capability should a general manipulation framework possess?
This paper categorizes foundation models trained using manipulation datasets as RFMs. RFMs significantly enhance the policy’s generalization. For instance, LEO (Huang et al. (2023a)) directly outputs target pose, while RoboFlamingo (Li et al. (2023c)) directly outputs delta pose. Directly outputting target pose is easier to train compared to delta pose, although delta pose better aligns with human behavior. Currently, RFMs are fine-tuned using pre-trained models, but this deprives them of self-exploratory learning. Designing an architecture for RFMs is crucial. Utilizing foundation models to aid reinforcement learning still holds many unexplored areas, such as using foundation models to assist robots in reinforcement learning within real-world environments.
Training RFMs simultaneously requires a large amount of training data. Current data collection methods include gathering data in simulation, using skill library to generate data, teleoperation, and imitation learning from human videos. While collecting data in simulation is cost-effective, it suffers from the sim-to-real gap. In order to generate data using skill library, it is necessary to construct an initial skill library. Teleoperation can address the sim-to-real gap but comes with high hardware costs and the issue of unnatural third-person demonstration data. Imitation learning from human videos involves mapping human trajectories from videos to robot space, leading to issues with trajectory accuracy. Which of these methods can effectively lead to general manipulation remains to be explored.
@@ 9.5 How to use internet-scale video data for RFMs?
Methods to obtain language annotations are divided into manual and automated captions. Manual captions are created by humans labeling video content. Automated captions include four types: (1) Automatic Speech Recognition (ASR), which converts audio in videos to text (Xue et al. (2022)). (2) Alt-text, which collects captions from HTML alt-text attributes in web images and videos, like descriptions, tags, and titles (Bain et al. (2021)). (3) Transfer, which starts with a set of image-caption pairs. Then, captions are matched to video clips with similar frames (Nagrani et al. (2022)). (4) Foundation Models, which use pre-trained models to get captions. For example, VLMs provide single-frame image captions, while LLMs filter out inconsistent captions across frames (Blattmann et al. (2023)). Owing to recent advancements in language annotation techniques, most widely used internet video datasets incorporate language annotations, such as InternVid (Wang et al. (2023c)), HD-VILA-100M (Xue et al. (2022)), YT-Temporal-180M (Zellers et al. (2021)), WTS-70M (Shvetsova et al. (2025)), HowTo100M (Miech et al. (2019)), WebVid-10M (Nan et al. (2024)), and VideoCC3M (Yan et al. (2022)).
@@
Transferring human actions from video to robot control is similar to cross-embodiment learning (Zakka et al. (2022)). Three key questions require focus: what information from video dataset can be used, how to extract these useful information and how to utilize the extracted information to enhance or train robotic foundation models. Regarding the first question, there are six main types of information to convert from video datasets: (1) Pose, such as capturing human hand poses and retargeting them to dexterous hand poses (Shaw et al. (2023); Qin et al. (2022)). (2) Affordance, including grasp locations on objects and post-grasp waypoints (Mendonca et al. (2023)). (3) Motion information, explicitly includes keypoints trajectories of objects and human hand during actions (Xiong et al. (2021); Yuan et al. (2024); Wen et al. (2023)) and implicitly includes using VQ-VAE (Van Den Oord et al. (2017)) to generate a codebook for latent delta action (Ye et al. (2024)). (4) Environment transition dynamic information, such as predicting hindsight images after completing the current action (Wu et al. (2023a); Cheang et al. (2024); Yang et al. (2023b)). (5) Semantic information, such as descriptions of current task steps (Wang et al. (2024a)) and task instruction (Jain et al. (2024)). (6) Spatial and texture information, such as MVP (Radosavovic et al. (2023)) suggests using masked autoencoding (He et al. (2022)) for improving visual reconstruction.
@@
Regarding the second question, the methods of language annotation have been introduced above. At the same time, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)). When labeling pose for videos, research on human pose detection and retargeting becomes crucial (Shaw et al. (2023); Qin et al. (2022); Li et al. (2024)). When adding various labels to the video dataset, different training objectives can be used to extract features from the video dataset, such as MAE (Radosavovic et al. (2023)), contrastive learning (Ma et al. (2022)), time contrastive learning (Ma et al. (2023a)), temporal-difference learning (Bhateja et al. (2023)), video prediction objective(Du et al. (2024)), affordance prediction objective(Mendonca et al. (2023)), video-language alignment objective (Nair et al. (2022)), action motion objective(Yuan et al. (2024)) or combinations of these objectives (Karamcheti et al. (2023); Zhou et al. (2021)).
@@
The current robotic foundation models primarily use two learning methods: imitation learning and reinforcement learning. Therefore, the discussion on the third issue focuses on leveraging prior knowledge from video datasets in these two methods. As for imitation learning, when the robotic foundation model outputs pose and the video dataset annoated with pose label, the video dataset can be directly used as training data for the robotic foundation model (Shaw et al. (2023); Qin et al. (2022)). When leveraging affordance information, motion information, environment transition dynamics information, semantic information, spatial and texture information, it is essential to employ GMM & CEM (Mendonca et al. (2023)), Inverse Dynamic Model (IDM) (Du et al. (2024); Ye et al. (2024); Wen et al. (2023)), and Decoder (Wang et al. (2023a); Xiao et al. (2022b); Cheang et al. (2024); Wu et al. (2023a)) to transform these information into actions. Compared to other types of information, using semantic information treats the video as task instruction rather than observation (Jain et al. (2024); Shah et al. (2023); Jang et al. (2022)). At the same time, semantic information can also be used to organize tasks into a hierarchy of skills (Wang et al. (2024a)).
@@ 9.6 How to design low-cost hardware for collecting large-scale real-world manipulation data?
For approaches capturing and mapping the pose of end-effectors, the common low-cost capturing devices include SpaceMouse (Liu et al. (2022a); Zhu et al. (2023)), cameras (Cheng et al. (2024); Fu et al. (2024); Iyer et al. (2024); Shaw et al. (2023); Li et al. (2019); Fang et al. (2020a)), VR controllers (De Pace et al. (2021); Nakanishi et al. (2020)) and IMU sensors (Chi et al. (2024); Li et al. (2020); Fang et al. (2017a, b)). The SpaceMouse based method passes the position and orientation of the SpaceMouse as action commands of end-effectors. This method is low-cost, easy operation, and easy implementation, but it is not suitable for dual-arm operations. In contrast, methods based on cameras and VR are well suited for bimanual teleoperation and VR offers the advantage of visual feedback compared to cameras. However, teleoperation methods based on cameras and VR heavily relies on the accuracy of pose estimation algorithms and often affected by occlusion (Fu et al. (2024); Pavlakos et al. (2024); Iyer et al. (2024); Cheng et al. (2024)). The main advantage of teleoperation devices based on IMU sensors lies in their wearability (Li et al. (2020); Chi et al. (2024); Wang et al. (2024b)). Due to this advantage, UMI (Chi et al. (2024)) and DexCap (Wang et al. (2024b)) develope wearable devices capable of in-the-wild teleoperation and offline data collection.
@@
As for teleoperation visual feedback, most of methods (Liu et al. (2022a); Zhu et al. (2023); Fu et al. (2024); Li et al. (2020); Zhao et al. (2023a); Wu et al. (2023c)) are use third side view that observe the robot task with the operator’s own eyes directly. However, this observation involves some visual errors. For example, there may be inaccuracies in the distance between the gripper and the object being manipulated. While for first-person view, due to wearing VR head (Cheng et al. (2024); Iyer et al. (2024); De Pace et al. (2021); Nakanishi et al. (2020)), it allows operators to perceive the robot’s surroundings immersively. However, long time to use VR headset can lead to fatigue.
@@
9.6.3 Summary.
Scene diversity primarily includes the diversity of scene layouts, such as floor plans and object placements, as well as the diversity of objects (Deitke et al. (2022)). Regarding the automatic generation of floor plans, current approaches are either rule-based (Deitke et al. (2022)) or learning-based (Yang et al. (2024)). However, they do not leverage the commonsense knowledge embedded in foundation models. Such knowledge can greatly enhance the naturalness of room layouts. For example, bedrooms are often connected to adjacent bathrooms through a door. Similarly, this commonsense knowledge applies to object layouts as well. For instance, knives are typically placed on the right side of plates, while forks are on the left (Ding et al. (2023)). Regarding the diversity of objects, although there are many existing 3D object assets (Chang et al. (2015); Deitke et al. (2023); Li et al. (2023a); Geng et al. (2023a); Xiang et al. (2020); Liu et al. (2022b); Calli et al. (2017)), their quantity is far from sufficient to cover the variety of real-world objects. As a result, generative models have gained popularity as a method for generating 3D objects. However, the performance of generative models is also limited by the shortage of current 3D object assets cannot cover objects in the real world. To address this issue, data cleaning techniques or manual supervision are needed to filter and select high-quality generated object assets. At the same time, generative models mainly focus on rigid and articulated objects and research on deformable objects remains insufficient (Sundaresan et al. (2022)).
As for task diversity, many studies currently utilize LLMs to generate instructions, followed by the use of skill libraries to create demonstrations (Wang et al. (2023b, d)). However, generating diverse task instructions using foundation models presents challenges in ensuring rationality. For example, integrating observations from specific scenarios to identify existing objects and their affordances is crucial. This approach helps make the instructions more feasible and actionable.

### v4

@@ 8.3 Sim-to-Real Gap Solutions
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (Zhang et al. (2024c)) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (Jiang et al. (2024a)) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
@@ 8.4 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (Mandlekar et al. (2023)) and DexMimicGen (Jiang et al. (2024c)) change the positions and orientations of objects, while CACTI (Mandi et al. (2022)) adds new, artificial objects to the scene. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Automated Generation methods. In the Real-to-Sim methods, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (Chen et al. (2024c)). Additionally, GRS (Zook et al. (2024)) utilizes VLM to extract 3D object meshes corresponding to real-world object from assert database based on RGB image. In the Automated Generation methods, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (Wang et al. (2023e)). When the output is scene code, it directly generates the corresponding scene (Wang et al. (2023c)). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (Wang et al. (2023e)). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@ 9 Discussion
In this survey, we aim to outline the opportunities brought by foundation models for general manipulation. We believe the potential of embedding foundation models into manipulation tasks as a viable path towards achieving general manipulation. However, the primary applications of LLMs, VFMs, VLMs, LMMs and VGMs focus only on certain aspects of general manipulation capability, such as reasoning, perception, multimodal understanding, and data generation. The current framework for RFMs demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring over a 99% success rate remains an unresolved concern. Therefore, this paper proposes a framework of robot learning for manipulation towards achieving general manipulation capability and detailing how foundation models can address challenges in each module of the framework. However, there are still many open questions in this survey. In this section, we delve into several open questions that we are particularly concerned about.
@@ 9.3 What kind of learning capability should a general manipulation framework possess?
This paper categorizes foundation models trained using manipulation datasets as RFMs. RFMs significantly enhance the policy’s generalization. For instance, LEO (Huang et al. (2023a)) directly outputs target pose, while RoboFlamingo (Li et al. (2023d)) directly outputs delta pose. Directly outputting target pose is easier to train compared to delta pose, although delta pose better aligns with human behavior. Currently, RFMs are fine-tuned using pre-trained models, but this deprives them of self-exploratory learning. Designing an architecture for RFMs is crucial. Utilizing foundation models to aid reinforcement learning still holds many unexplored areas, such as using foundation models to assist robots in reinforcement learning within real-world environments.
Training RFMs simultaneously requires a large amount of training data. Current data collection methods include gathering data in simulation, using skill library to generate data, teleoperation, and imitation learning from human videos. While collecting data in simulation is cost-effective, it suffers from the sim-to-real gap. In order to generate data using skill library, it is necessary to construct an initial skill library. Teleoperation can address the sim-to-real gap but comes with high hardware costs. Imitation learning from human videos involves mapping human trajectories from videos to robot space, leading to issues with retargeting. Which of these methods can effectively lead to general manipulation remains to be explored.
@@ 9.5 How to use internet-scale video data for RFMs?
Methods to obtain language annotations are divided into manual and automated captions. Manual captions are created by humans labeling video content. Automated captions include four types: (1) Automatic Speech Recognition (ASR), which converts audio in videos to text (Xue et al. (2022)). (2) Alt-text, which collects captions from HTML alt-text attributes in web images and videos, like descriptions, tags, and titles (Bain et al. (2021)). (3) Transfer, which starts with a set of image-caption pairs. Then, captions are matched to video clips with similar frames (Nagrani et al. (2022)). (4) Foundation Models, which use pre-trained models to get captions. For example, VLMs provide single-frame image captions, while LLMs filter out inconsistent captions across frames (Blattmann et al. (2023)). Owing to recent advancements in language annotation techniques, most widely used internet video datasets incorporate language annotations, such as InternVid (Wang et al. (2023d)), HD-VILA-100M (Xue et al. (2022)), YT-Temporal-180M (Zellers et al. (2021)), WTS-70M (Shvetsova et al. (2025)), HowTo100M (Miech et al. (2019)), WebVid-10M (Nan et al. (2024)), and VideoCC3M (Yan et al. (2022)).
@@
Transferring human actions from video to robot control is similar to cross-embodiment learning (Zakka et al. (2022)). Three key questions require focus: what information from video dataset can be used, how to extract these useful information and how to utilize the extracted information to enhance or train robotic foundation models. Regarding the first question, there are six main types of information to convert from video datasets: (1) Pose, such as capturing human hand poses and retargeting them to dexterous hand poses (Shaw et al. (2023); Qin et al. (2022)). (2) Affordance, including grasp locations on objects and post-grasp waypoints (Mendonca et al. (2023)). (3) Motion information, explicitly includes keypoints trajectories of objects and human hand during actions (Xiong et al. (2021); Yuan et al. (2024); Wen et al. (2023b)) and implicitly includes using VQ-VAE (Van Den Oord et al. (2017)) to generate a codebook for latent delta action (Ye et al. (2024)). (4) Environment transition dynamic information, such as predicting hindsight images after completing the current action (Wu et al. (2023a); Cheang et al. (2024); Yang et al. (2023b)). (5) Semantic information, such as descriptions of current task steps (Wang et al. (2024a)) and task instruction (Jain et al. (2024)). (6) Spatial and texture information, such as MVP (Radosavovic et al. (2023)) suggests using masked autoencoding (He et al. (2022)) for improving visual reconstruction.
@@
Regarding the second question, the methods of language annotation have been introduced above. At the same time, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023b)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)). When labeling pose for videos, research on human pose detection and retargeting becomes crucial (Shaw et al. (2023); Qin et al. (2022); Li et al. (2024)). When adding various labels to the video dataset, different training objectives can be used to extract features from the video dataset, such as MAE (Radosavovic et al. (2023)), contrastive learning (Ma et al. (2022)), time contrastive learning (Ma et al. (2023a)), temporal-difference learning (Bhateja et al. (2023)), video prediction objective(Du et al. (2024)), affordance prediction objective(Mendonca et al. (2023)), video-language alignment objective (Nair et al. (2022)), action motion objective(Yuan et al. (2024)) or combinations of these objectives (Karamcheti et al. (2023); Zhou et al. (2021)).
@@
The current robotic foundation models primarily use two learning methods: imitation learning and reinforcement learning. Therefore, the discussion on the third issue focuses on leveraging prior knowledge from video datasets in these two methods. As for imitation learning, when the robotic foundation model outputs pose and the video dataset annoated with pose label, the video dataset can be directly used as training data for the robotic foundation model (Shaw et al. (2023); Qin et al. (2022)). When leveraging affordance information, motion information, environment transition dynamics information, semantic information, spatial and texture information, it is essential to employ GMM & CEM (Mendonca et al. (2023)), Inverse Dynamic Model (IDM) (Du et al. (2024); Ye et al. (2024); Wen et al. (2023b)), and Decoder (Wang et al. (2023a); Xiao et al. (2022b); Cheang et al. (2024); Wu et al. (2023a)) to transform these information into actions. Compared to other types of information, using semantic information treats the video as task instruction rather than observation (Jain et al. (2024); Shah et al. (2023); Jang et al. (2022)). At the same time, semantic information can also be used to organize tasks into a hierarchy of skills (Wang et al. (2024a)).
@@ 9.6 How to design low-cost hardware for collecting large-scale real-world manipulation data?
For approaches capturing and mapping the pose of end-effectors, the common low-cost capturing devices include SpaceMouse (Liu et al. (2022a); Zhu et al. (2023b)), cameras (Cheng et al. (2024); Fu et al. (2024); Iyer et al. (2024); Shaw et al. (2023); Li et al. (2019); Fang et al. (2020a)), VR controllers (De Pace et al. (2021); Nakanishi et al. (2020)) and IMU sensors (Chi et al. (2024); Li et al. (2020); Fang et al. (2017a, b)). The SpaceMouse based method passes the position and orientation of the SpaceMouse as action commands of end-effectors. This method is low-cost, easy operation, and easy implementation, but it is not suitable for dual-arm operations. In contrast, methods based on cameras and VR are well suited for bimanual teleoperation and VR offers the advantage of visual feedback compared to cameras. However, teleoperation methods based on cameras and VR heavily relies on the accuracy of pose estimation algorithms and often affected by occlusion (Fu et al. (2024); Pavlakos et al. (2024); Iyer et al. (2024); Cheng et al. (2024)). The main advantage of teleoperation devices based on IMU sensors lies in their wearability (Li et al. (2020); Chi et al. (2024); Wang et al. (2024b)). Due to this advantage, UMI (Chi et al. (2024)) and DexCap (Wang et al. (2024b)) develope wearable devices capable of in-the-wild teleoperation and offline data collection.
@@
As for teleoperation visual feedback, most of methods (Liu et al. (2022a); Zhu et al. (2023b); Fu et al. (2024); Li et al. (2020); Zhao et al. (2023a); Wu et al. (2023c)) are use third side view that observe the robot task with the operator’s own eyes directly. However, this observation involves some visual errors. For example, there may be inaccuracies in the distance between the gripper and the object being manipulated. While for first-person view, due to wearing VR head (Cheng et al. (2024); Iyer et al. (2024); De Pace et al. (2021); Nakanishi et al. (2020)), it allows operators to perceive the robot’s surroundings immersively. However, long time to use VR headset can lead to fatigue.
9.6.3 Summary.
Scene diversity primarily includes the diversity of scene layouts, such as floor plans and object placements, as well as the diversity of objects (Deitke et al. (2022)). Regarding the automatic generation of floor plans, current approaches are either rule-based (Deitke et al. (2022)) or learning-based (Yang et al. (2024b)). However, they do not leverage the commonsense knowledge embedded in foundation models. Such knowledge can greatly enhance the naturalness of room layouts. For example, bedrooms are often connected to adjacent bathrooms through a door. Similarly, this commonsense knowledge applies to object layouts as well. For instance, knives are typically placed on the right side of plates, while forks are on the left (Ding et al. (2023)). Regarding the diversity of objects, although there are many existing 3D object assets (Chang et al. (2015); Deitke et al. (2023); Li et al. (2023a); Geng et al. (2023a); Xiang et al. (2020); Liu et al. (2022b); Calli et al. (2017)), their quantity is far from sufficient to cover the variety of real-world objects. As a result, generative models have gained popularity as a method for generating 3D objects. However, the performance of generative models is also limited by the shortage of current 3D object assets cannot cover objects in the real world. To address this issue, data cleaning techniques or manual supervision are needed to filter and select high-quality generated object assets. At the same time, generative models mainly focus on rigid and articulated objects and research on deformable objects remains insufficient (Sundaresan et al. (2022)).
@@ 9.7 How to collect large-scale diversity and realism manipulation data in simulation?
As for task diversity, many studies currently utilize LLMs to generate instructions, followed by the use of skill libraries to create demonstrations (Wang et al. (2023c, e)). However, generating diverse task instructions using foundation models presents challenges in ensuring rationality. For example, integrating observations from specific scenarios to identify existing objects and their affordances is crucial. This approach helps make the instructions more feasible and actionable.